### v1

@@ 2.2 Instrumental Variables
Consider an endogenous binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$. Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$. Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction. Let $Y=Y(D,Z)$ be the realized outcome. As before, let $X$ denote covariates. We make the following assumptions.
@@
    First stage: $e(X)\equiv\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$ almost surely.
We also make one of the following two nested monotonicity assumptions.
Assumption 2.3 (Strong monotonicity).
$\mathbb{P}(D(1)\geq D(0)\mid X)=1$ almost surely.
Assumption 2.4 (Weak monotonicity).
There exists a subset of the support of $X$ such that $\mathbb{P}(D(1)\geq D(0)\mid X)=1$ on it and $\mathbb{P}(D(1)\leq D(0)\mid X)=1$ on its complement.
The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995). In addition to the assumptions on the instrument, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e. let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, where $\text{supp}(\cdot)$ is used to denote the support. Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\mathbbm{1}(X=x_{1}),\ldots,Z\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector. The estimand in Angrist and Imbens (1995) is the following 2SLS estimand:
$\beta_{\text{2SLS}}\equiv\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector. This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the specific representation in Proposition 2.2 follows Słoczyński (2020).
@@
Let Assumptions 2.2 and 2.4 hold, and let $X$ be discrete with finite support. Then
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[e(X)(1-e(X))\cdot\mathbb{P}\left(D(1)\neq D(0)\mid X\right)^{2}\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]\right]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}.$
This means that we can write $\beta_{\text{2SLS}}$ as
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=\mathbbm{1}(D(1)\neq D(0))$ is the population of compliers/defiers, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and $a(X)=e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, a nonnegative weight function. Note that $\beta_{\text{2SLS}}=\text{LATE}\equiv\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0)]$ if and only if $a(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)\neq D(0)$.
The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create additional instruments by interacting the original instrument with covariates (cf. Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is constructed to obtain $\beta_{\text{2SLS}}$ above. A more practically relevant estimand is the “just-identified” IV estimand,
$\beta_{\text{IV}}\equiv\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$. To introduce one of the representations of $\beta_{\text{IV}}$ below, define
$c(X)\equiv\operatorname{sign}\Big{(}\mathbb{P}\left[D(1)\geq D(0)\mid X\right]-\mathbb{P}\left[D(1)\leq D(0)\mid X\right]\Big{)},$
where $\operatorname{sign}(\cdot)$ is the sign function. We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 2.2.
@@
$e(X)$ is linear in $X$.
Under the instrument validity assumption, the rich covariates assumption, and either monotonicity assumption, Słoczyński (2020) obtains the following representations for the “just-identified” IV estimand.
@@
Let Assumptions 2.2, 2.4, and 2.5 hold. Then
$\beta_{\text{IV}}=\frac{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]]}{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)]}.$
If Assumptions 2.2, 2.3, and 2.5 hold instead, then
$\beta_{\text{IV}}=\frac{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)]}.$
It follows that we can write $\beta_{\text{IV}}$ as
$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}$
under either monotonicity assumption. Under weak monotonicity, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and possibly negative weights $a(X)=c(X)\cdot e(X)(1-e(X))$. Under strong monotonicity, $W_{0}=\mathbbm{1}(D(1)>D(0))$, $w_{0}(X)=\mathbb{P}(D(1)>D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and nonnegative weights $a(X)=e(X)(1-e(X))\geq 0$.
@@ 2.3 Difference-in-Differences
Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$. We assume units are untreated prior to period $G\in\mathcal{G}=\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter. We assume no units are treated in the first time period. This may include a group that remains untreated throughout, for which $G=+\infty$. Thus, $D_{t}=\mathbbm{1}(G\leq t)$. The panel is balanced, that is, no group appears or disappears over time.
The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators. Using partitioned regression results, it is defined as
$\beta_{\text{TWFE}}\equiv\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right]},$
@@
    $\mathcal{G}\equiv\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
@@
    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\mathcal{G}$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of the two-way fixed effects estimand as a weighted average.
@@
Let Assumption 2.6 hold. Then
$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}}.$
We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$. This period variable denotes the time period and we use it to define $(Y(1),Y(0))\equiv(Y_{P}(1),Y_{P}(0))$, $Y=Y_{P}$, and $D=D_{P}$, which are potential outcomes, the realized outcome, and treatment at a random time period, respectively.
@@
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G,P)\in\{0,1\}$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$, and the weight function $a(X)=1-\mathbb{P}(D=1\mid G)-\mathbb{P}(D=1\mid P)+\mathbb{P}(D=1)$ is not generally nonnegative. [^2] Here, $\tau_{0}(X)$ is what Callaway and Sant’Anna (2021) call “the group-time average treatment effect.” A nonnegative weight function can be obtained under the assumption that group-level average treatment effects are constant over time. This property was described in Goodman-Bacon (2021, Section 3.1.1) and a representation of the two-way fixed effects estimand is given in his equation (16). The following proposition yields a simpler (but equivalent) expression for the weights in our setting.
@@
Let Assumption 2.6 hold and let $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]},$
where $a_{B}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.

### v3

@@ 5.2   Instrumental Variables
Now consider a binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$. Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$. Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction. Let $Y=Y(D,Z)$ be the realized outcome. As before, let $X$ denote covariates. We make the following assumptions.
@@
    First stage: $\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$;
  4. 4.
    Strong monotonicity: $\mathbb{P}(D(1)\geq D(0)\mid X)=1$.
The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995). In addition to Assumption 5.1, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e., let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$. Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\cdot\mathbbm{1}(X=x_{1}),\ldots,Z\cdot\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector. The estimand in Angrist and Imbens (1995) is given by
$\beta_{\text{2SLS}}\coloneqq\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector. This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the representation in Proposition 5.1 follows from Słoczyński (2020).
@@
Suppose Assumption 5.1 holds. Suppose $X$ is discrete with finite support. Then
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[|\operatorname{cov}(D,Z\mid X)|\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]\mid D(1)>D(0)\right]}{\mathbb{E}[|\operatorname{cov}(D,Z\mid X)|\mid D(1)>D(0)]}.$
Thus, $\beta_{\text{2SLS}}$ satisfies the representation in (1.1) with $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and $W_{0}=\mathbbm{1}(D(1)>D(0))$. Note that $\beta_{\text{2SLS}}=\text{LATE}\coloneqq\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0)]$ if and only if $a_{\text{2SLS}}(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)>D(0)$.
The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create multiple interacted instruments (cf. Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is defined and used to obtain $\beta_{\text{2SLS}}$ above. A more practically relevant estimand is the “noninteracted” IV estimand,
$\beta_{\text{IV}}\coloneqq\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$. We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 5.1.
@@
$\mathbb{P}(Z=1\mid X)$ is linear in $X$.
Under the instrument validity assumption and the rich covariates assumption, Słoczyński (2020) obtains the following representation of the “noninteracted” IV estimand.
@@
Suppose Assumptions 5.1 and 5.2 hold. Then
$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]\mid D(1)>D(0)]}{\mathbb{E}[\operatorname{var}(Z\mid X)\mid D(1)>D(0)]}.$
It follows that $\beta_{\text{IV}}$ is a weighted estimand satisfying (1.1) with weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$, CATE function $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and where the average is again taken over the complier subpopulation, i.e., $W_{0}=\mathbbm{1}(D(1)>D(0))$.
Causal Representation and Internal Validity of 2SLS
First consider the estimand $\beta_{\text{2SLS}}$, which can be characterized as $\mu(a_{\text{2SLS}},\tau_{0})$. Since $a_{\text{2SLS}}(X)\geq 0$, there exists a subpopulation of $\{D(1)>D(0)\}$ such that $\beta_{\text{2SLS}}$ is an average treatment effect over that subpopulation. The maximum size of that subpopulation is given by
$\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{2SLS}}(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a_{\text{2SLS}}(x)}=\frac{\mathbb{E}[|\operatorname{cov}(D,Z\mid X)|\mid D(1)>D(0)]}{\sup_{x\in\text{supp}(X\mid D(1)>D(0))}|\operatorname{cov}(D,Z\mid X=x)|}.$
The maximum value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is obtained when $|\operatorname{cov}(D,Z\mid X)|$ does not depend on $X$. This occurs, for example, when the instrument and the fraction of units for which $D(1)>D(0)$ are independent of $X$. In this case, we have that $\beta_{\text{2SLS}}=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0)]$, the average treatment effect for compliers.
Under the representation in Proposition 5.2, the IV estimand has the same $W_{0}$, but has $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ instead. Here, $a_{\text{IV}}(X)\geq 0$ and
$\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{IV}}(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a_{\text{IV}}(x)}=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\mid D(1)>D(0)]}{\sup_{x\in\text{supp}(X\mid D(1)>D(0))}\operatorname{var}(Z\mid X=x)}.$
The internal validity of the IV estimand is maximized when $\operatorname{var}(Z\mid X)$ is constant, which occurs when $Z$ is independent of $X$. In this case, $\beta_{\text{IV}}$ equals LATE. The quantities $\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ and $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ are not ranked uniformly in the distributions of $(D(1),D(0),X,Z)$ as there are data-generating processes that make each of these two quantities larger than the other. For example, if $\operatorname{var}(Z\mid X)$ is constant but $\mathbb{P}(D(1)>D(0)\mid X)$ is not, then $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})<\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$. This scenario is plausible if $Z$ is randomly assigned and $X$ is a vector of pre-assignment characteristics. This inequality is reversed if $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|$ is constant but $\mathbb{P}(D(1)>D(0)\mid X)$ is not. They are equally representative when $\mathbb{P}(D(1)>D(0)\mid X)$ is constant. In this case, the estimands are equal, so this is not unexpected.
@@ 5.3   Difference-in-Differences
Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment by $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$. We assume units are untreated prior to period $G\in\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter. We assume no units are treated in the first time period. This may include a group that remains untreated throughout, for which $G=+\infty$. Thus, $D_{t}=\mathbbm{1}(G\leq t)$. The panel is balanced, that is, no group appears or disappears over time.
The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators. By partitioned regression results, the coefficient on treatment indicator is
$\beta_{\text{TWFE}}\coloneqq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]\big{/}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right],$
@@
    $\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
@@
    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\text{supp}(G)$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of $\beta_{\text{TWFE}}$ as a weighted average.
@@
Suppose Assumption 5.3 holds. Then
$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}a_{\text{TWFE}}(G,t)\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}a_{\text{TWFE}}(G,t)\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}},$
where $a_{\text{TWFE}}(g,t)=1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G=g]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$. This period variable denotes the time period and we use it to define $(Y(1),Y(0),Y,D)\coloneqq(Y_{P}(1),Y_{P}(0),Y_{P},D_{P})$, which are potential outcomes, the realized outcome, and treatment at random period $P$, respectively.
@@
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{\text{TWFE}}(X)\cdot\tau_{0}(X)\mid D=1]}{\mathbb{E}[a_{\text{TWFE}}(X)\mid D=1]},$
[…]
@@
Suppose Assumption 5.3 holds and that $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]\mid D=1]}{\mathbb{E}[a_{\text{TWFE,H}}(G)\mid D=1]},$
where $a_{\text{TWFE,H}}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.